OpenStudy (anonymous):

If the expected relationship you are investigating is likely to be of the form y = 2.0/(x1/3) and you plot a graph of log y vs log x what would you expect the value of the slope to be , if the expected relationship is observed?

OpenStudy (anonymous):

i think form of the question is y = 2.0/(x^1/3) so by taking log log y=log2-1/3 (logx) so this is the curve for which you have to find slope so differentiate both side wrt to x you would find (1/y)dy/dx =-(1/3x) so slope dy/dx =-y/3x

OpenStudy (anonymous):

thanks a lot much appreciated!! hope to help you where I can

OpenStudy (anonymous):

you are welcome @marcus91